Supervisor: “We sign two new customers out of every seven people with whom we speak. Our goal is to sign 120 new customers. How many people will we have to speak with to meet our goal?”
Representative: “We will have to speak with __________ people.“

Answer by jim_thompson5910(35256)

(2/7)*x = 120

2x = 120*7

2x = 840

x = 840/2

x = 420


Supervisor: “We sign two new customers out of every seven people with whom we speak. Our goal is to sign 120 new customers. How many people will we have to speak with to meet our goal?”

Representative: “We will have to speak with _____420_____ people.“